Lucy Lightfoot stands with one foot on one bathroom scale and her other foot on a second bathroom scale. Each scale reads 325 N. What was Lucy's weight?
1
Expert's answer
2019-05-29T09:41:23-0400

We can find Lucy's weight from the Newton's Second Law of Motion:


"\sum F_y = 0,""N_1 + N_2 - W_{Lucy} = 0,"

here, "W_{Lucy}" is Lucy's weight, directed downward, "N_1, N_2" are the normal forces (or the scale readings) directed upward.

Then, from this equation we can find Lucy's weight:


"W_{Lucy} = N_1 + N_2 = 325 N + 325 N = 650 N."

Answer:

"W_{Lucy} = 650 N."